OpenStudy (anonymous):

. Essay; show all work. A simple dartboard has three areas… the main board has a radius of 9 inches, there is a circle with a radius of 5 inches, and the bullseye has a radius of 3 inches. What is the probability of a random dart landing inside the bullseye? Round to the nearest thousandth. (Hint: Use A= (pie r2

OpenStudy (anonymous):

area of the bulls eye divided by area of the dart board

OpenStudy (anonymous):

\[\frac{\pi\times 3^2}{\pi \times 9^2}\] in other words multiply, cancel

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

\[\frac{9\pi}{81\pi}=\frac{1}{9}\]

OpenStudy (anonymous):

1/9 sorry

OpenStudy (anonymous):

what would the probabilitty be then

OpenStudy (anonymous):

1 out of 9

OpenStudy (anonymous):